Assume the following information for company SDA on 31st December 2018, before
the preparation of the adjusted trial-balance:
Machinery (€) 38,000
Wages payable 1,300
Accumulated depreciation: machinery 6,000
Bank deposits 12,000
Unearned revenue 2,100
Prepaid insurance 3,600
Accounts receivable 6,500
Advances from customers 1,800
Revenue 13,000
Inventory 3,000
Long-term loans 16,000
Accounts payable 4,500
Retained earnings 13,000
Rent expense 9,600
Share Capital 15,000

Required

Prepare adjusting entries by taking into consideration the following
information:
a. The accrued interest on long-term loans for the fiscal year 2018 is calculated at an
interest rate of 6% on the balance of the loan at this date.
b. During December, unearned revenues of €1,500 were earned.
c. Depreciation charges on machinery for fiscal year 2018 amount to €3,160.
d. The accrued interest on deposits in December is €170.
e. On 10th October 2018, SDA prepaid €4,800 for six months’ rent and debited the
account Rent expense for the total amount of €4,800.
f. An amount of €600 was recognized as revenue of fiscal year 2018. This amount
corresponds to services that will be provided in the next fiscal year 2019.
g. Prepaid insurance expired at 31st December 2018.

Record the closing journal entries for the determination of the firm’s net
profit or loss assuming that the firm’s tax rate is 20% and the annual result of its
operations (profit or loss) is carried forward

Prepare the post-closing trial balance.
Provide the required calculations. Assume that the fiscal year end date of SDA is on
31st December 2018.
Answer #1

Answer:

Date Debit (€) Credit (€) 960 Dec.31 960 1,500 Dec.31 1,500 3,160 Dec.31 3,160 Adjusting General Journal Account Name Interes

Rent expense paid in advance on 10 Oct 2018 = 4,800 but wrongly debited to prepaid rent.

Hence in prepaid rent 4,800 balance is there.

Upto December 2018 expired prepaid rent is = 4,800/6 × 2.67 month= 2,136

For October only for 20 days i.e. 20/30= .67 month is considered for calculation.

So, correct rent expense balance = 4,800 - 2,136 = 2,664

And correct prepaid rent balance = 4,800 - 2,136 = 2,664

Debit Credit 13,900 170 14070 14,653 Closing General Journal Date Account Names and explanations Revenue earned Interest reve

POST CLOSING TRIAL BALANCE Account Name Debit Credit Interest payable € 960 Unearned revenue 1,200 Accumulated depreciation 9